OpenStudy (anonymous):

Find the requested function value f(x)=6x+9, g(x)=-5x^2+9x-9 Find (g o f)(7)

OpenStudy (alfie):

You'll have to compose the two functions. Begin by doing the function composition between these two functions.

OpenStudy (anonymous):

(g o f) = -5(6x+9)^2 + 9(6x+9) + 9 (g o f)(7) = -5(51)^2 + 9(51) + 9 - ill let you compute this!

OpenStudy (anonymous):

basically its a composite function - you replace every x in g by f which is 6x+9

OpenStudy (anonymous):

- which is my first line - then you replace x by 7 fot calcuate (g o f) (7)